Consider the line y=6/5x+3 . The x is in the middle of the fraction 6/5x.
What is the slope of a line perpendicular to this line?
What is the slope of a line parallel to this line?

Answer by tutorcecilia(2152)


Perpendicular lines have negative reciprocal slopes:
Slope #1 = 6/5
Slope #2 = -5/6
.
Parallel lines have the same slope:
Slope #1 = 6/5
Slope #2 = 6/5